OpenStudy (anonymous):

I need help real bad its algebra 1 semester 2 anyone please :)

OpenStudy (anonymous):

how can i do this http://prntscr.com/36hzme

OpenStudy (solomonzelman):

y ≥ x -1 is a one unit shift down from x=y In which to f the choices is the line graphed y=x-1 ? In which f these two choices where the line is y=x-1 is the shaded space above the y?

OpenStudy (solomonzelman):

graph the line "y ≥ x -1 " on you paper just like you would graph "y=x-1" then shade everything above the line